OpenStudy (anonymous):

the value of a 1950 nickel is 7/4 the value of a 1945 nickel, the total value to the coins is 55.00, what is the value of each coin?

OpenStudy (lgbasallote):

what is the unit of that 55.00? dollars? cents?

OpenStudy (anonymous):

dollars

OpenStudy (lgbasallote):

let x = value of 1945 nickel so value of 1950 nickel is 7/4 x it says the total is $55.00 so \[x + \frac 74 x = 55\] convert to common denominators \[\implies \frac 44 x + \frac 74 x = 55\] combine like terms \[\implies \frac{11}4 x = 55\] multiply both sides by 4 \[\implies 11x = 55 \times 4\] simplify \[\implies 11x = 220\] now divide both sides by 11 \[x = 220 \div 11\] simplify \[x = 20\] are you sure the unit is dollars? because according to this thing, the value of a 1945 nickel is 20 dollars

OpenStudy (anonymous):

that is exactly the answer that the problem came up with but i couldn't figure out how it came up with that answer thank you for your help it explained it for me. :)

OpenStudy (lgbasallote):

welcome :D

OpenStudy (lgbasallote):

20 dollars for a nickel is still weird though